OpenStudy (anonymous):

Solve using the elimination method.show your work. If the system has no solution or an infinite number of solutions, state this. 11x+10y=77 , 9x+y=63

OpenStudy (hba):

11x+10y=77 10y=77-11x y=77-11x/10

OpenStudy (hba):

now put it in the other equation

OpenStudy (hba):

9(77-11x/10)+y=63

OpenStudy (hba):

y=63-9(77-11x/10)

OpenStudy (hba):

y=63+(693-99x/10)

OpenStudy (hba):

nw take lcm

OpenStudy (hba):

y=630-693+99x/10

OpenStudy (mathteacher1729):

\[ \begin{align} 11x&+10y&=77 \\ 9x&+y&=63 \end{align} \] If you want to eliminate a variable you can multiply the bottom equation by -10 \[ \begin{align} 11x &+10y &=77 \\ -90x&-10y &=-630 \end{align} \] Now eliminating y should be relatively easy to do. Can you solve the rest of the problem from here?

OpenStudy (hba):

oops i thought he meant substitution

OpenStudy (mathteacher1729):

First sentence: "Solve using the elimination method." ;-p